Determine the volume occupied by 10 kg water at a pressure of 10 MPa and a temperature of (a)5"C, (b)200 "C, (c) 400 "C, (d)800 "C, (e) 1500 "C, and (f)-10 "C. Ans. (a)0.00996 m3 (b)0.0115 m3 (c) 0.2641 m3 (d)0.4859 m3 (e)0.8182 m3 (f)0.01089 m3
2
1
Expert's answer
2020-08-09T08:43:58-0400

The volume of water at the given temperature and pressure is found by using impressibility factor (Z).

Actual volume (V)="\frac{mZRT}{P}"

Z is obtained from impressibility chart by using reduced pressure(Pr) and reduced temperature(Pr).

Tr=T/Tc,

Pr=T/Pc

where Tc and Pc are values of temp and pressure at critical point.

the calculation is given in tabular form for all questions

.